there are 9 books 55% of 9 books are out,how many are left? solve this problem using a proportion show all steps.and explain
Answer by nerdybill(6958)

there are 9 books 55% of 9 books are out,how many are left? solve this problem using a proportion show all steps.and explain
.
since 55% are out
100-55 = 45% are left
.
Let x = books left
then
x/9 = 45/100 (this is your proportion)
cross multiplying:
100x = 9*45
100x = 405
x = 4.05
since you can't have fractional book, you have to round down:
there are 4 books left